Dx/y^2+z^2-x^2=dy/-2xy=dz/-2xy
1
Expert's answer
2021-09-01T12:58:16-0400

From the equation "\frac{dy}{-2xy}=\frac{dz}{-2xy}" we can deduce that "dy=dz" and that the function "f_1=y-z" is the first integral of the given system.

Consider now the first equation "\frac{dx}{y^2+z^2-x^2}=\frac{dy}{-2xy}".

(1) "0=2xydx+dy(y^2+z^2-x^2)=y^2 \left( \frac{2xydx-x^2dy}{y^2} +(1+\frac{z^2}{y^2})dy\right)"

But

(2) "\frac{2xydx-x^2dy}{y^2}=d(x^2/y)"

(3) "(1+\frac{z^2}{y^2})dy=(1+\frac{(y-f_1)^2}{y^2})dy=(2-\frac{2f_1}{y}+\frac{f_1^2}{y^2})dy="

"=d(2y-2f_1\log |y|-f_1^2/y)"

Substituting eq.(2),(3) into eq.(1), we obtain "d(x^2/y+2y-2f_1\log |y|-f_1^2/y)=0"

Therefore the function "f_2=x^2/y+2y-2(y-z)\log |y|-(y-z)^2/y" is also the integral of the given system.

"df_1\wedge df_2=df_1\wedge d(x^2/y)+df_1\wedge d (2y-2f_1\log |y|-f_1^2/y)="

"df_1\wedge d(x^2/y)+(1+\frac{z^2}{y^2})df_1\wedge dy="

"y^{-2}df_1\wedge (2xydx+(y^2+z^2-x^2)dy)\ne 0" inside the open dense set

"U=\{(x,y,z)\in\mathbb{R}^3:xy(y^2+z^2-x^2)\ne0\}".

Therefore, the functions "f_1" and "f_2" are functionally independent in "U", and the complete integral for the given system in the domain "U" can be given by the formula "\Phi(f_1,f_2)" with an arbitrary smooth function "\Phi".

Answer. "\Phi(y-z, x^2y^{-1}+2y-2(y-z)\log |y|-y^{-1}(y-z)^2)" is the complete integral.